# On the Emergence of Position Bias in Transformers

### v2

@@ 5 Experiments
Following Reddy (2024), we adopt the following information retrieval task: The model is trained to predict the label $y_{\text{query}}$ of a target $x_{\text{query}}$ using the cross-entropy loss, given an alternating sequence of $n$ items and $n$ labels: $x_{1},y_{1},\dots,x_{n},y_{n},x_{\text{query}}$. The sequence is embedded in $d$ dimensions. Each $x_{i}$ is sampled from a Gaussian mixture model with $K$ classes, and $y_{i}$ is the corresponding class label assigned prior to training from the total $L$ labels ($L\leq K$). The burstiness $B$ is the number of occurrences of $x_{i}$ from a particular class in an input sequence. Importantly, at least one item in the context belongs to the same class as the query. To control position bias in the training data, $x_{\text{query}}$ can either be explicitly assigned to the class of a specific $x_{i}$, introducing position-dependent bias in the data, or randomly assigned to the class of any $x_{i}$, simulating a scenario without position bias in the data.
@@
Figure 2: Position bias arising solely from the architectural design of the attention mechanism, with no positional bias in the training data. $a$ vs. $b$ denotes the gap for the case $[\bm{a},b]-[\bm{b},a]$, where bar magnitude indicates gap size, positive indicates bias toward earlier position, and negative indicates bias toward later position. Deeper attention amplifies the bias toward earlier tokens, regardless of the PE used. Furthermore, decay mask introduce stronger distance-based decay effects that increase focus on recent tokens than RoPE.
@@ 5.1 The Effects of Depth and Relative PEs
To investigate the position bias arising solely from the architectural design of the attention mechanism, we use training sequences without positional bias, where the position of $x_{i}$ sharing the same class as $x_{\text{query}}$ is uniformly random in $\{1,2,\ldots,n\}$. To evaluate the position bias in the trained model, we construct test sequences of the form $[\bm{a},b]$. Here, the bolded term $\bm{a}$ explicitly marks the correct position, ensuring $y_{a}$ matches $y_{\text{query}}$, while position $b$ serves as a baseline. In these sequences, $x_{a}$ and $x_{b}$ are identical vectors, allowing us to control for the influence of semantic information on the model’s retrieval accuracy. We then measure the retrieval accuracy gap between pairs of sequences where the content at positions $a$ and $b$ is identical, but the correct position varies. This gap, defined as $[\bm{a},b]-[\bm{b},a]$, quantifies the model’s positional preference independent of semantic information. To perform this evaluation, we construct three pairs of test sets, each containing $10,000$ sequences: [first, middle] vs. [middle, first], [first, last] vs. [last, first], and [middle, last] vs. [last, middle]. Here “first” (position 1), “middle” (position $n/2$), and “last” (position $n$) denote fixed positions within a sequence.
@@
Figure 2 shows the average results over five runs, where $a$ vs. $b$ denotes the gap $[\bm{a},b]-[\bm{b},a]$. The magnitude of each bar represents the size of the performance gap, and the sign of each bar reflects the direction of the bias: a positive sign indicates a bias toward earlier positions, while a negative sign indicates a bias toward later positions. We highlight several key observations. First, increasing model depth consistently amplifies the bias toward earlier parts of the sequence, regardless of the PE used. Also note that the performance gap between the middle and last positions is notably smaller than that involving the first position. This aligns with our theory, which suggests that as the model focuses more on the initial part of the sequence, information near the sequence’s end becomes less distinguishable, consistent with the patterns observed in Barbero et al. (2024a). Furthermore, both the decay mask and RoPE introduce distance-based decay effects that reduce the bias toward the beginning induced by the causal mask and increase the focus on recent tokens. However, the decay effect induced by the decay mask is substantially more pronounced than that by RoPE, as predicted by our theory.
@@ 5.2 Can Causal Mask Induce Usable Positional Information?
Next, we empirically examine how the causal mask leverages positional information. Kazemnejad et al. (2023) hypothesized that without PE (No PE), the causal mask can implicitly simulate absolute or relative PE through specific weight matrix configurations. To test this hypothesis, we train models on sequences with positional bias at either the beginning or the end. Specifically, in the training data, $x_{\text{query}}$ is assigned to the class of $x_{1}$ or $x_{n}$ with equal probability. We then evaluate two types of attention masks: no mask ($\mathcal{G}$ is complete) and causal, and three types of PEs: No PE (Kazemnejad et al., 2023), absolute sinusoidal PE (sin PE) (Vaswani et al., 2017), and relative PE using RoPE. For evaluation, we construct six types of test sets as described in Section 5.1, each with $10,000$ sequences.
@@
It is worth noting that in Figure 3, we observe the “lost-in-the-middle” phenomenon (Liu et al., 2024), where information retrieval accuracy follows a U-shape relative to the position of the answer, with performance at the beginning of the sequence slightly better than at the end. More experimental results under different types of positional bias in the training data can be found in Section K.1. Notably, this phenomenon does not occur when the training data lacks positional bias (Figure 2) or contains other types of positional bias considered (Section K.1). This suggests that specific types of positional bias in the training data also play a role in how the model learns to prioritize positions within a sequence.
@@ 6 Conclusion
In this paper, we study position bias in transformers through a probabilistic and graph-theoretic lens, developing a theoretical framework that quantifies how positional information influences context construction across multi-layer attention. Our analysis reveals two key findings about position bias in transformers: the causal mask’s inherent bias toward earlier tokens, as deeper layers increasingly attend to these positions through iterative attention, and the interplay between causal masking and relative positional encodings, which results in a nuanced, non-monotonic balance between distance-based decay effects and the cumulative influence of earlier positions. These findings open several promising directions for future work. One potential direction is leveraging these insights to design bias-free transformers, mitigating positional biases to improve model robustness and generalization capabilities. Alternatively, our framework can also inform the strategic exploitation of positional bias in specific applications, such as emphasizing early positions for text summarization or prioritizing recent interactions in recommendation systems. Another important direction is to extend our analysis to other transformer components—such as value projections and MLPs—which may interact with attention in nontrivial ways, as suggested by analyses based on vector norms (Kobayashi et al., 2020). By deepening our understanding of how architectural choices in transformers shape positional dependencies, our work provides a foundation for designing attention mechanisms with predictable and task-aligned positional properties.
@@ I.1 Results
As defined in Section 5, the input data distribution is modulated by tuning various parameters. In addition to the parameters described in the main text, for the Gaussian mixture with $K$ classes, each class $k$ is defined by a $d$-dimensional vector $\mu_{k}$ whose components are sampled $i.i.d.$ from a normal distribution with mean zero and variance $1/d$. Then the value of $x_{i}$ is given by $\frac{\mu_{k}+\gamma\eta}{\sqrt{1+\gamma^{2}}}$, where $\eta$ is drawn from the same distribution as the $\mu_{k}$’s and $\gamma$ sets the within-class variability. Each class is assigned to one of $L$ labels ($L\leq K$). The contents of the labels are drawn prior to training from the same distribution as the $\mu_{k}$’s.
@@
K.1 The role of training data on positional bias
@@
Interestingly, when comparing this behavior to the case shown in Figure 3, we note that the “lost-in-the-middle” phenomenon only emerges when the training sequences are biased toward both the beginning and the end. This suggests that specific types of positional bias in the training data play a crucial role in shaping how the model learns to process and prioritize positions within a sequence.

### v3

@@ 5 Experiments
Following Reddy (2024), we adopt the following information retrieval task: The model is trained to predict the label $y_{\text{query}}$ of a target $x_{\text{query}}$ using the cross-entropy loss, given an alternating sequence of $n$ items and $n$ labels: $x_{1},y_{1},\dots,x_{n},y_{n},x_{\text{query}}$. The sequence is embedded in $d$ dimensions. Each $x_{i}$ is sampled from a Gaussian mixture model with $K$ classes (see Appendix L for results under a fixed-vocabulary setup), and $y_{i}$ is the corresponding class label assigned prior to training from the total $L$ labels ($L\leq K$). The burstiness $B$ is the number of occurrences of $x_{i}$ from a particular class in an input sequence. Importantly, at least one item in the context belongs to the same class as the query. To control position bias in the training data, $x_{\text{query}}$ can either be explicitly assigned to the class of a specific $x_{i}$, introducing position-dependent bias in the data, or randomly assigned to the class of any $x_{i}$, simulating a scenario without position bias in the data.
@@
Figure 2: Position bias arising solely from the architectural design of the attention mechanism, with no positional bias in the training data. $a$ vs. $b$ denotes the gap for the case $[\bm{a},b]-[\bm{b},a]$, where bar magnitude indicates gap size, positive indicates bias toward earlier position, and negative indicates bias toward later position. Deeper attention amplifies the bias toward earlier tokens, regardless of the PE used. Furthermore, decay mask introduce stronger distance-based decay effects that increase focus on recent tokens than RoPE. Residual connections affect positional bias in a nontrivial, depth- and PE-dependent way (see Section K.2 for more details).
@@ 5.1 The Effects of Depth and Relative PEs
To investigate the position bias arising solely from the architectural design of the attention mechanism, we use training sequences without positional bias, where the position of $x_{i}$ sharing the same class as $x_{\text{query}}$ is uniformly random in $\{1,2,\ldots,n\}$. We evaluate two architectural variants: one without residual connections, and one with standard residual connections applied after each attention layer. The no-residual variant serves as a theoretical probe, isolating the contribution of attention alone to positional bias. In contrast, the residual variant reflects standard architectural design and enables us to examine how residual connections, in addition to attention, influence the emergence and amplification of positional bias.
@@
To evaluate the position bias in the trained model, we construct test sequences of the form $[\bm{a},b]$. Here, the bolded term $\bm{a}$ explicitly marks the correct position, ensuring $y_{a}$ matches $y_{\text{query}}$, while position $b$ serves as a baseline. In these sequences, $x_{a}$ and $x_{b}$ are identical vectors, allowing us to control for the influence of semantic information on the model’s retrieval accuracy. We then measure the retrieval accuracy gap between pairs of sequences where the content at positions $a$ and $b$ is identical, but the correct position varies. This gap, defined as $[\bm{a},b]-[\bm{b},a]$, quantifies the model’s positional preference independent of semantic information. To perform this evaluation, we construct three pairs of test sets, each containing $10,000$ sequences: [first, middle] vs. [middle, first], [first, last] vs. [last, first], and [middle, last] vs. [last, middle]. Here “first” (position 1), “middle” (position $n/2$), and “last” (position $n$) denote fixed positions within a sequence.
@@
Figure 2 shows the average results over five runs, where $a$ vs. $b$ denotes the gap $[\bm{a},b]-[\bm{b},a]$. The magnitude of each bar represents the size of the performance gap, and the sign of each bar reflects the direction of the bias: a positive sign indicates a bias toward earlier positions, while a negative sign indicates a bias toward later positions. We highlight several key observations. First, without residual connections, increasing model depth consistently amplifies the bias toward earlier parts of the sequence, regardless of the PE used (see Section K.1 for additional results). Also note that the performance gap between the middle and last positions is notably smaller than that involving the first position. This aligns with our theory, which suggests that as the model focuses more on the initial part of the sequence, information near the sequence’s end becomes less distinguishable, consistent with the patterns observed in Barbero et al. (2024a). Furthermore, both the decay mask and RoPE introduce distance-based decay effects that reduce the bias toward the beginning induced by the causal mask and increase the focus on recent tokens. However, the decay effect induced by the decay mask is substantially more pronounced than that by RoPE, as predicted by our theory.
Effect of residual connections on position bias
Residual connections appear to interact with positional bias in a nontrivial way. While position bias still emerges in models with residual connections, its relationship with depth becomes non-monotonic and can be either amplified or reduced, depending on the positional encoding used and the depth regime. We provide a more detailed analysis in Section K.2.
@@ 5.2 Can Causal Mask Induce Usable Positional Information?
Next, we empirically examine how the causal mask leverages positional information. Kazemnejad et al. (2023) hypothesized that without PE (No PE), the causal mask can implicitly simulate absolute or relative PE through specific weight matrix configurations. To test this hypothesis, we train models with residual connections on sequences with positional bias at either the beginning or the end. Specifically, in the training data, $x_{\text{query}}$ is assigned to the class of $x_{1}$ or $x_{n}$ with equal probability. We then evaluate two types of attention masks: no mask ($\mathcal{G}$ is complete) and causal, and three types of PEs: No PE (Kazemnejad et al., 2023), absolute sinusoidal PE (sin PE) (Vaswani et al., 2017), and relative PE using RoPE. For evaluation, we construct six types of test sets as described in Section 5.1, each with $10,000$ sequences.
@@
It is worth noting that in Figure 3, we observe the “lost-in-the-middle” phenomenon (Liu et al., 2024), where information retrieval accuracy follows a U-shape relative to the position of the answer, with performance at the beginning of the sequence slightly better than at the end. More experimental results under different types of positional bias in the training data can be found in Section K.3. Notably, this phenomenon does not occur when the training data lacks positional bias (Figure 2) [^2] With the sole exception occurring in the deep-layer regime with RoPE; see Section K.2. or contains other types of positional bias considered (Section K.3). This suggests that specific types of positional bias in the training data also play a role in how the model learns to prioritize positions within a sequence.
An additional observation from the fixed-vocabulary setting (see Appendix L for details) is that the geometry of the token embeddings can significantly influence task difficulty. This geometric structure may also affect the emergence and nature of positional bias, as it shapes how the model learns to distinguish and attend to tokens.
@@ 6 Conclusion
In this paper, we study position bias in transformers through a probabilistic and graph-theoretic lens, developing a theoretical framework that quantifies how positional information influences context construction across multi-layer attention. Our analysis reveals two key findings about position bias in transformers: the causal mask’s inherent bias toward earlier tokens, as deeper layers increasingly attend to these positions through iterative attention, and the interplay between causal masking and relative positional encodings, which results in a nuanced, non-monotonic balance between distance-based decay effects and the cumulative influence of earlier positions. These findings open several promising directions for future work. One potential direction is leveraging these insights to design bias-free transformers, mitigating positional biases to improve model robustness and generalization capabilities. Alternatively, our framework can also inform the strategic exploitation of positional bias in specific applications, such as emphasizing early positions for text summarization or prioritizing recent interactions in recommendation systems. Another important direction is to extend our analysis to other transformer components—such as residual connections, value projections, and MLPs—which may interact with attention in nontrivial ways, as suggested by our results in Section 5.1 and analyses based on vector norms (Kobayashi et al., 2020). By deepening our understanding of how architectural choices in transformers shape positional dependencies, our work provides a foundation for designing attention mechanisms with predictable and task-aligned positional properties.
@@ I.1 Results
As defined in Section 5, the input data distribution is modulated by tuning various parameters. In addition to the parameters described in the main text, for the Gaussian mixture with $K$ classes, each class $k$ is defined by a $d$-dimensional vector $\mu_{k}$ whose components are sampled $i.i.d.$ from a normal distribution with mean zero and variance $1/d$. Then the value of $x_{i}$ is given by $\mu_{k}+\epsilon\eta_{i}/\sqrt{1+\epsilon^{2}}$, where $\eta_{i}$ is drawn from the same distribution as the $\mu_{k}$’s and $\epsilon$ sets the within-class variability. Each class is assigned to one of $L$ labels ($L\leq K$). The contents of the labels are drawn prior to training from the same distribution as the $\mu_{k}$’s.
@@
K.1 The effect of depth and relative PEs
Here, we present additional experimental results on the effect of attention depth on positional bias. Since models without residual connections become increasingly difficult to train as depth increases, we limit our experiments to depths of 2, 4, and 6. The results are shown in Figure 4. We observe that deeper attention amplifies the bias toward earlier tokens, regardless of the PE used. Furthermore, both decay mask and RoPE introduce distance-based decay effects that increase focus on recent tokens, but the one induced by decay mask is stronger.
Figure 4: Results with pure attention layers: position bias arising solely from the architectural design of the attention mechanism, with no positional bias in the training data.
K.2 The effect of residual connections on position bias
In this section, we present additional experimental results on the effect of residual connections on the positional bias exhibited by the model. The results are shown in Figure 5. As discussed in Section 5.1, we observe that with residual connections, the relationship between model depth and positional bias becomes non-monotonic, and the strength of the bias can be either amplified or reduced depending on the positional encoding used and the depth regime. These findings suggest that residual connections modulate the accumulation of positional effects across layers in a more complex manner than with pure attention alone. It would be interesting to further explore and characterize how residuals influence the emergence and amplification of positional bias.
Figure 5: Results with residual connections: position bias arising solely from the architectural design of the attention mechanism, with no positional bias in the training data.
K.3 The role of training data on positional bias
@@
Interestingly, when comparing this behavior to the case shown in Figure 3, we note that the “lost-in-the-middle” phenomenon only emerges when the training sequences are biased toward both the beginning and the end (the only other exception occurs in the deep-layer regime with RoPE, despite the absence of positional bias in the training data; see Section K.2). This suggests that specific types of positional bias in the training data could also play a crucial role in shaping how the model learns to process and prioritize positions within a sequence.
@@
Appendix L Additional Results under a Fixed-Vocabulary Setting
In addition to the Gaussian mixture setting discussed in the main text, we also present results for a fixed-vocabulary setting: we define a fixed set of $K$ class vectors $\{v_{1},\dots,v_{K}\}$, and assign each input $x_{i}=v_{c(i)}$, where $c(i)\in\{1,\dots,K\}$ is the class index. Each class vector $v_{k}$ is then defined by $v_{k}=(\mu_{k}+\lambda\omega)/\sqrt{1+\lambda^{2}}$, where $\mu_{k}$ and $\omega$ are $d$-dimensional vectors whose components are sampled $i.i.d.$ from a normal distribution with mean zero and variance $1/d$. Note that the shared component $\eta$ makes $v_{k}$ not completely orthogonal to each other, resembling token embeddings in practice (Cai et al., 2021; Godey et al., 2024; Ethayarajh, 2019; Gao et al., 2019; Ait-Saada & Nadif, 2023). $\lambda$ is a scalar that controls the strength of anisotropy, where we set $\lambda=0.75$.
The corresponding results are shown in Figure 9-Figure 13, respectively. Most of the findings presented in the main text continue to hold in this setting. Notably, even with residual connections, increasing depth consistently amplifies positional bias, as shown in Figure 10.
One key difference between the two settings is that, compared to the orthogonal embeddings used in the Gaussian mixture setting of Reddy (2024), the anisotropic embeddings in our setup appear to make the task significantly more challenging. We note that in preliminary experiments under the fixed-vocabulary setting with $\lambda=0$, the results closely resemble those from the Gaussian mixture setup, suggesting that embedding geometry plays a critical role in task difficulty. It is possible that residual connections behave differently when the task becomes more difficult. Future work could further investigate how the geometry of the input embeddings influences both task difficulty and the emergence of positional bias.
Figure 9: Results with pure attention layers under the fixed-vocabulary setting: position bias arising solely from the architectural design of the attention mechanism, with no positional bias in the training data.
Figure 10: Results with residual connections under the fixed-vocabulary setting: position bias arising solely from the architectural design of the attention mechanism, with no positional bias in the training data.
(a) Position bias exhibited by the model when trained on data biased toward the first and last positions.
(b) Position bias exhibited by the model when trained on data biased toward the first position.
(c) Position bias exhibited by the model when trained on data biased toward the middle position.
(d) Position bias exhibited by the model when trained on data biased toward the last position.
Figure 11: Results under fixed-vocabulary setting
Figure 12: Fixed-vocabulary setting: example of the emergence of attention sinks in our experimental setting. In particular, the sequences used for training and inference are all free of position bias.
Figure 13: Attention sinks under the fixed-vocabulary setting.